# TC-Light: Temporally Coherent Generative Rendering for Realistic World Transfer

### v2

@@ 4.2 Comparison with SOTA
Quantitative and qualitative comparisons with state-of-the-art methods are reported in Tab. 2 and Fig. 3, respectively. The result indicates that per-frame relighting (IC-Light*) follows prompts well and produces physically plausible illumination, but the adapted illumination suffers from severe flicker, as shown in columns (a) and (b) of Fig. 3. IC-Light would even randomly hallucinate non-existent objects in textureless regions (cf. column (c) of Fig. 3), further degrading consistency. Extending IC-Light* with VidToMe li2024vidtome yields modest gains in temporal coherence but dramatically lowers computation cost for long videos, so we adopt it as our primary baseline. Slicedit cohen2024slicedit significantly suppresses flicker and hallucinations, yet its computation overhead exceeds that of IC-Light*. Besides, its output remains overly biased by the original appearance of the source video. As a result, it produces unnatural relighting in many cases, as shown in column (a) of Fig. 3.
We also evaluated the T2V-model-based video editing approach wang2025videodirector and concurrent video relighting techniques zhou2025light ; fang2025relightvid . Unfortunately, they all failed on long clips due to OOM errors caused by high computation resource demands. For the same reason, Cosmos-Transfer1 alhaija2025cosmos can only operate in single-modality mode under GPU constraints, yet still requires over 30 GB GPU memory and more than 30 minutes per clip. Moreover, on video with high dynamics, it suffers from more severe blur and loss of details, as shown in columns (a) and (b) of Fig. 3. These failures are likely because Cosmos-Transfer1 is limited to the data domain of its training data, which contains less varied, moderately dynamic videos.
In contrast, our TC-Light first enables physically plausible relighting on long videos with high dynamics, while outperforming all baselines in temporal consistency and preference rate by a large margin, as shown in Tab. 2. The light version adds only 2.4 minutes and 2.7 GB of VRAM overhead compared to the VidToMe baseline, while faithfully preserving object identity, albedo, and adherence to text prompts, as shown in Fig. 3. Incorporating our decayed multi-axis denoising further enhances temporal coherence, with a modest trade-off in efficiency and quality. Limited by page, we provide additional visualization and performance of different scenarios types in the Appendix.
@@ 4.3 Ablation
Table 4: Ablation over Unique Video Tensor (UVT). Here, %$\mathrm{Cmpr}$ is the compression rate after applying UVT on the source video. The subscripts “f” and “f+d” indicate that, besides color cues, the UVT representation incorporates optical flow cues and both flow and depth cues, respectively.
Scene %$\mathrm{Cmpr}_{f}$ $\mathrm{SSIM}_{f}$ $\mathrm{PSNR}_{f}$ $\mathrm{LPIPS}_{f}$$\mathrm{Cmpr}_{f+d}$ $\mathrm{SSIM}_{f+d}$ $\mathrm{PSNR}_{f+d}$ $\mathrm{LPIPS}_{f+d}$
@@
Figure 4: Ablation on main module components. The experiment is conducted on one sequence of the InteriorNet InteriorNet18 subset, where the text prompt is "This video showcases a modern interior space, which is dimly lit". The baseline here denotes VidToMe li2024vidtome in Tab. 2.
This section analyzes the contribution of each component in our model. The first stage optimization, as shown in Tab. 3 and Fig. 4, markedly boosts consistency by aligning cross-frame exposure. The 6-7th rows of Tab. 3 also illustrate that, initializing UVT optimization from the first-stage results converges more efficiently than directly optimizing UVT for the same overall epochs. The second stage optimization, as shown in Tab. 3, further reinforces temporal coherence. Tab. 4 confirms that UVT can compress the source video with near-zero loss, which underpins our design in Sec. 3.3.2. Using UVT as the second-stage target not only boosts consistency but also cuts computational overhead. Additionally, replacing a hard mask with a soft mask consistently improves both Warp-SSIM and CLIP-T metrics, demonstrating its importance. Incorporating the depth cues alongside UVT yields a more compact representation (also in Tab. 4), which aids illumination alignment and release computation burden. In contrast, instance segmentation masks provide no clear benefit and are thus omitted from the final implementation.
For the diffusion module, multi-axis denoising notably enhances temporal consistency. However, it tends to inherit appearance distribution from the source video, causing drift from the target prompt and sometimes unnatural lighting, as shown in Fig. 4. The introduced AIN and weight decay mitigate these issues, achieving a promising balance between consistency and faithful prompt alignment.
@@ 4.4 Limitation and Discussion
Despite achieving impressive results, our method is still limited by its base models. For instance, the current version of IC-Light zhang2025scaling still struggles to relight hard shadows or make large modifications to low-light images. Similarly, since IC-Light is pretrained on 512 resolution and fine-tuned on 1024 resolution, our model struggles to preserve image details if the resolution is lower than 512. Besides, since the optimization process relies on the optical flow estimation model, artifacts sometimes occurs in textureless areas where flow becomes unreliable. Furthermore, the temporal consistency loss has the tendency to smooth the texture of flickering areas, and therefore might sacrifice some details. Though the proposed decayed multiaxis denoising alleviates the problem, developing a temporally more consistent and computationally more efficient denoising strategy is desired in future work.
@@ 5 Conclusion
In summary, we present TC-Light, a one-shot-tuned framework that delivers temporally consistent and physically plausible relighting on long, highly dynamic videos. The optimization-based illumination alignment provides a new paradigm for video relighting. Central to our approach is the Unique Video Tensor—an explicit, canonical, and differentiable video representation that enables highly efficient optimization. Over the established long video relighting benchmark, TC-Light achieves state-of-the-art performance in both consistency and efficiency, making it particularly well suited for sim2real and real2real data scaling in embodied AI.
@@ Appendix A Additional Experimental Results
Fig. 5 presents additional visualizations of our relighting results across a diverse range of scenarios. Whether under nighttime or daytime conditions, in outdoor or indoor environments, or from aerial or ground-level viewpoints, the proposed TC-Light method consistently produces temporally coherent and physically plausible illumination edits, demonstrating strong generalization capabilities. Fig. 6 provides qualitative comparisons against state-of-the-art methods across additional scenarios. As shown, our model effectively adheres to textual instructions while generating relighting results that are both natural and temporally consistent.
We also provide corresponding quantitative evaluations on synthetic and real-world scenarios. As reported in Tab. 5, performance on real-world scenes consistently exceeds that on synthetic ones. This discrepancy likely arises from the training data of the video model Cosmos-Transfer1 alhaija2025cosmos and the foundational image model IC-Light zhang2025scaling , which are biased towards realistic scenes. Furthermore, the higher resolution and richer textures of real-world data mitigate hallucinations in textureless regions and help better preserve the intrinsic details of source frames for IC-Light. Such attributes are particularly critical for the consistency of methods with comparatively limited temporal modeling, namely, IC-Light* and VidToMe, which exhibit substantially higher Motion-S and WarpSSIM metrics on real-world videos than on synthetic ones. In contrast, our approach attains state-of-the-art temporal consistency across both scenario types while maintaining a favorable balance with prompt adherence.
@@ Appendix B Details of Assets
License N/A CC-BY Custom [^1] https://waymo.com/open/terms/ CC BY- NC-SA 4.0 CC BY- NC-SA 4.0 Apache -2.0 Custom [^2] https://interiornet.org/ CC0 1.0 N/A
In Tab. 6, we summarize the license and resolution for each subset. All source videos are resized and center-cropped to their designated resolutions. Considering the computation source limitation, we sample videos that contain over 600 frames to around 300 frames. The DRONE subset includes three clips captured using our DJI Mini4 Pro and two additional clips obtained from DroneStock [^3] https://dronestock.com/, which are released under the CC0 1.0 License. For AgiBot Digital World contributors2025agibotworld , where the robot’s head moves in coordination with its body while performing tasks, relighting is performed from the head-mounted camera view. For each scene of DROID khazatsky2024droid , we apply relighting to both the static side camera and the dynamic left wrist camera views. For Waymo sun2020scalability and NavSim Dauner2024NEURIPS , relighting is conducted using the front-facing camera view.
@@ Appendix C Additional Implementation Details
For competing methods, we adopt the hyperparameters from their official implementations for VideoDirector wang2025videodirector , Light-A-Video zhou2025light , RelightVid fang2025relightvid , and Cosmos-T1 alhaija2025cosmos . We replace base models of VidToMe li2024vidtome and Slicedit cohen2024slicedit with IC-Light zhang2025scaling , and therefore we align their classifier-free guidance scale and diffusion sampling steps with those in zhang2025scaling . Additionally, we set VidToMe’s local and global token-merging ratios to 0.6 and 0.5, respectively, mirroring the setting of our approach. For Slicedit, we adjust the weighting factor $\gamma$ in Eq. (2) from the default 0.2 to 0.05 to better balance temporal coherence and instruction adherence. All other hyperparameters remain at their default values. The modified VidToMe serves as the baseline of our model design.
During implementation, each $\kappa(x,y,t)$ comprises three components: (1) a per-pixel flow ID, (2) a quantized RGB color, and, optionally, (3) a world-frame voxel coordinate. For (1), flow IDs are derived from the optical flow estimated by the state-of-the-art MemFlow method dong2024memflow and the binary mask obtained by thresholding the soft mask in Eq. (6) of the main paper (values > 0.5 are set to 1; otherwise 0). In the initial frame, pixels receive unique flow IDs from 00 to $HW-1$, where $H$ and $W$ denote image height and width. In subsequent frames, a pixel inherits the flow ID of its predecessor if connected by an unmasked flow; otherwise, it is assigned a new ID. This injects motion priors into the UVT representation. For (2), we quantize RGB values to 7 bits, ensuring that all pixels sharing the same UVT element differ by less than $2/255$ in any channel. This constraint mitigates erroneous flows that escape the mask and reinforces representation in regions exhibiting view-dependent effects. For (3), when per-frame depth maps are available, they are reprojected into a point cloud using the camera intrinsics and extrinsics to determine world-frame coordinates. This point cloud is then voxelized at a specified voxel size, and each pixel’s voxel coordinate is appended to $\kappa(x,y,t)$, yielding a more compact representation of static regions. For the CARLA dosovitskiy2017carla and InteriorNet InteriorNet18 datasets, voxel sizes are set to 0.05 m and 0.02 m, respectively. Notably, dynamic objects at different timestamps may spatially overlap in 3D, but they remain distinguishable by their flow IDs and quantized RGB colors. Consequently, each object at each timestep is represented by a distinct set of UVT elements, while the $\mathcal{L}_{1}$ temporal consistency loss preserves object identity across frames.
@@ Appendix D User Study
We conducted an online user study with 78 anonymous participants, evaluating 19 randomly selected video–text pairs from our datasets. The compared methods were IC-Light* zhang2025scaling , VidToMe li2024vidtome , Slicedit cohen2024slicedit , Cosmos-Transfer1 alhaija2025cosmos , Ours-light, and Ours-full. A screenshot of the questionnaire interface is shown in Fig. 7. For each question, methods were anonymized and relighted videos were presented in random order; participants selected the two most preferred results. In compliance with the NeurIPS Code of Ethics, each participant received a compensation of \$0.70. Besides, we ensured that all collected data remained confidential and was not disclosed to any institutions or individuals.
Since each video spanned 10–20 seconds, completing the questionnaire took on average 13.5 minutes. Submissions requiring less than four minutes were deemed unreliable and excluded, yielding 65 valid responses. Fig. 8 reports the frequency with which each method was chosen among the top two. Our full model achieved the highest preference rate, while the light variant ranked second. Although IC-Light* and VidToMe follow instructions well (cf. Tab. 2 of the main paper), their inferior temporal consistency make them much less preferred by users. Finally, we computed Bradley–Terry preference scores bradley1952rank as a comprehensive metric of user preference, as presented in the Tab. 2 of the main paper.
@@
Appendix E Social Impact

### v3

@@ 4.2 Comparison with SOTA
Table 3: Comparison with existing methods on the Virtual KITTI 2 dataset (cabon2020vkitti2). The symbol definition aligns with Tab.˜2. Experiments are conducted on 40G A100. The best and the second best of each metric are separately highlighted in red and blue.
Method SSIM↑ LPIPS↓ Motion-S↑ Warp-SSIM↑ Time(s)↓ VRAM(G)↓
IC-Light* zhang2025scaling 0.5102 0.4470 95.23 68.13 1770 10.25
VidToMe li2024vidtome 0.5359 0.4262 95.95 71.33 444 6.96
Slicedit cohen2024slicedit 0.5080 0.4237 96.91 80.74 2346 17.68
VideoDirector wang2025videodirector OOM OOM OOM OOM OOM OOM
Light-A-Video zhou2025light OOM OOM OOM OOM OOM OOM
RelightVid fang2025relightvid OOM OOM OOM OOM OOM OOM
Cosmos-T1 alhaija2025cosmos 0.4833 0.4841 97.81 84.35 3314 34.83
Ours-light 0.5855 0.4026 98.51 90.94 580 15.21
Ours-full 0.5910 0.3971 98.62 92.38 1002 15.21
Quantitative and qualitative comparisons with state-of-the-art methods are reported in Tab.˜2, Tab.˜3, and Fig.˜3. The result indicates that per-frame relighting (IC-Light*) follows prompts well and produces physically plausible illumination, but the adapted illumination suffers from severe flicker, as shown in columns (a) and (b) of Fig.˜3. IC-Light would even randomly hallucinate non-existent objects in textureless regions (cf. column (c) of Fig.˜3), further degrading consistency. Extending IC-Light* with VidToMe li2024vidtome yields modest gains in temporal coherence but dramatically lowers computation cost for long videos, so we adopt it as our primary baseline. Slicedit cohen2024slicedit significantly suppresses flicker and hallucinations, yet its computation overhead exceeds that of IC-Light*. Besides, its output remains overly biased by the original appearance of the source video. As a result, it produces unnatural relighting in many cases, as shown in column (a) of Fig.˜3.
We also evaluated the T2V-model-based video editing approach wang2025videodirector and concurrent video relighting techniques zhou2025light; fang2025relightvid. Unfortunately, they all failed on long clips due to OOM errors caused by high computation resource demands. For the same reason, Cosmos-Transfer1 alhaija2025cosmos can only operate in single-modality mode under GPU constraints, yet still requires over 30 GB GPU memory and more than 30 minutes per clip. Moreover, on video with high dynamics, it suffers from more severe blur and loss of details, as shown in columns (a) and (b) of Fig.˜3. These failures are likely because Cosmos-Transfer1 is limited to the data domain of its training data, which contains less varied, moderately dynamic videos.
In contrast, our TC-Light first enables physically plausible relighting on long videos with high dynamics, while outperforming all baselines in temporal consistency and preference rate by a large margin, as shown in Tab.˜2. Tab.˜3 also demonstrates that our model outperforms all baselines in perceptual and structural similarity with ground truth relighting results. Considering computation cost, the light version adds only 2.4 minutes and 2.7 GB of VRAM overhead compared to the VidToMe baseline (cf. Tab.˜2), while faithfully preserving object identity, albedo, and adherence to text prompts, as shown in Fig.˜3. Incorporating our decayed multi-axis denoising further enhances temporal coherence, with a modest trade-off in efficiency and quality. Limited by page, we provide additional visualization and performance of different scenarios types in the Appendix. And the video demos and comparison can be found on our project page.
@@ 4.3 Ablation
Table 5: Ablation over Unique Video Tensor (UVT). Here, $\mathrm{Cmpr}$ is the compression rate after applying UVT on the source video. The subscripts “f” and “f+d” indicate that, besides color cues, the UVT representation incorporates optical flow cues and both flow and depth cues, respectively.
Scene $\mathrm{Cmpr}_{f}$ $\mathrm{SSIM}_{f}$ $\mathrm{PSNR}_{f}$ $\mathrm{LPIPS}_{f}$ $\mathrm{Cmpr}_{f+d}$ $\mathrm{SSIM}_{f+d}$ $\mathrm{PSNR}_{f+d}$ $\mathrm{LPIPS}_{f+d}$
@@
Figure 4: Ablation on main module components. The experiment is conducted on one sequence of the InteriorNet InteriorNet18 subset, where the text prompt is "This video showcases a modern interior space, which is dimly lit". The baseline here denotes VidToMe li2024vidtome in Tab.˜2.
This section analyzes the contribution of each component in our model. The first stage optimization, as shown in Tab.˜4 and Fig.˜4, markedly boosts consistency by aligning cross-frame exposure. The 6-7th rows of Tab.˜4 also illustrate that, initializing UVT optimization from the first-stage results converges more efficiently than directly optimizing UVT for the same overall epochs. The second stage optimization, as shown in Tab.˜4, further reinforces temporal coherence. Tab.˜5 confirms that UVT can compress the source video with near-zero loss, which underpins our design in Sec.˜3.3.2. Using UVT as the second-stage target not only boosts consistency but also cuts computational overhead. Additionally, replacing a hard mask with a soft mask consistently improves both Warp-SSIM and CLIP-T metrics, demonstrating its importance. Incorporating the depth cues alongside UVT yields a more compact representation (also in Tab.˜5), which aids illumination alignment and release computation burden. In contrast, instance segmentation masks provide no clear benefit and are thus omitted from the final implementation.
For the diffusion module, multi-axis denoising notably enhances temporal consistency. However, it tends to inherit appearance distribution from the source video, causing drift from the target prompt and sometimes unnatural lighting, as shown in Fig.˜4. The introduced AIN and weight decay mitigate these issues, achieving a promising balance between consistency and faithful prompt alignment.
@@ 4.4 Limitation and Discussion
Despite achieving impressive results, our method is still limited by its base models. For instance, the current version of IC-Light zhang2025scaling still struggles to relight hard shadows or make large modifications to low-light images. Similarly, since IC-Light is pretrained on 512 resolution and fine-tuned on 1024 resolution, our model struggles to preserve image details if the resolution is lower than 512. For instance, downscaling the NavSim subset from 960×536 to 480×264 causes Warp-SSIM to drop from the average value 90.46 to 88.36, although CLIP-T remains at 0.304. Besides, since the optimization process relies on the optical flow estimation model, artifacts sometimes occur in textureless areas or under very fast motion, where the flow becomes unreliable. For instance, when downsampling videos from the NavSim subset by 4 times to simulate very fast motion, the CLIP-T and Motion Smoothness fluctuate less than 2%, while Warp-SSIM declines by  5%. Furthermore, the temporal consistency loss has the tendency to smooth the texture of flickering areas, and therefore might sacrifice some details. It can be observed by comparing our model with IC-Light* in part (c) of Fig.˜3. Though the proposed decayed multiaxis denoising alleviates the problem, developing a temporally more consistent and computationally more efficient denoising strategy is desired in future work.
@@ 5 Conclusion
In summary, we present TC-Light, a one-shot-tuned framework that delivers temporally consistent and physically plausible relighting on long, highly dynamic videos. The optimization-based illumination alignment provides a new paradigm for video relighting. Central to our approach is the Unique Video Tensor—an explicit, canonical, and differentiable video representation that enables highly efficient optimization. Over the established long video relighting benchmark, TC-Light achieves state-of-the-art performance in both consistency and efficiency, endowing it with value and potential for broader application areas such as sim2real and real-world video scaling in embodied AI training and validation pipelines.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (No. 62320106010, No. U21B2042) and the Guangdong Provincial Foshan Joint Funds (No. 2024A1515110065).
Appendix
@@ Appendix A Additional Experimental Results
Fig.˜5 presents additional visualizations of our relighting results across a diverse range of scenarios. Whether under nighttime or daytime conditions, in outdoor or indoor environments, or from aerial or ground-level viewpoints, the proposed TC-Light method consistently produces temporally coherent and physically plausible illumination edits, demonstrating strong generalization capabilities. It is also worth noticing that the top row demonstrates our model’s ability to handle spatially varying lighting. In the middle three images, a white car drives from left to right. Initially, it is illuminated by orange street lamps, reflecting an orange hue. As it moves right, its rear remains orange-lit, while the front becomes blue due to a nearby advertising screen. Eventually, the car is fully bathed in blue light. This dynamic lighting response indicates our model can correctly handle spatially varying light. Fig.˜6 provides qualitative comparisons against state-of-the-art methods across additional scenarios. As shown, our model effectively adheres to textual instructions while generating relighting results that are both natural and temporally consistent.
We also provide corresponding quantitative evaluations on synthetic and real-world scenarios. As reported in Tab.˜6, performance on real-world scenes consistently exceeds that on synthetic ones. This discrepancy likely arises from the training data of the video model Cosmos-Transfer1 alhaija2025cosmos and the foundational image model IC-Light zhang2025scaling, which are biased towards realistic scenes. Furthermore, the higher resolution and richer textures of real-world data mitigate hallucinations in textureless regions and help better preserve the intrinsic details of source frames for IC-Light. Such attributes are particularly critical for the consistency of methods with comparatively limited temporal modeling, namely, IC-Light* and VidToMe, which exhibit substantially higher Motion-S and WarpSSIM metrics on real-world videos than on synthetic ones. In contrast, our approach attains state-of-the-art temporal consistency across both scenario types while maintaining a favorable balance with prompt adherence.
@@ Appendix B Details of Assets
License N/A $C$-BY Custom [^1] https://waymo.com/open/terms/
$C$ BY-
NC-SA 4.0
$C$ BY-
NC-SA 4.0
Apache
-2.0
Custom [^2] https://interiornet.org/ CC0 1.0 N/A
In Tab.˜7, we summarize the license and resolution for each subset. All source videos are resized and center-cropped to their designated resolutions. Considering the computation source limitation, we keep all frames if the sequence length is shorter than 300, and randomly sample 300 consecutive frames otherwise. Statistics are provided in Table 1 of the main paper. The DRONE subset includes three clips captured using our DJI Mini4 Pro and two additional clips obtained from DroneStock [^3] https://dronestock.com/, which are released under the CC0 1.0 License. For AgiBot Digital World contributors2025agibotworld, where the robot’s head moves in coordination with its body while performing tasks, relighting is performed from the head-mounted camera view. For each scene of DROID khazatsky2024droid, we apply relighting to both the static side camera and the dynamic left wrist camera views. For Waymo sun2020scalability and NavSim Dauner2024NEURIPS, relighting is conducted using the front-facing camera view. Domain balance is maintained between synthetic and real environments (25 vs. 28 videos), also balanced within sub-domains: autonomous driving (12 synthetic, 10 real), robotic manipulation (8 synthetic, 12 real), indoor navigation (5 synthetic, 6 real). The aerial subset is excluded from balance due to limited long dynamic drone videos in the simulation environment and serves mainly for generalization validation
@@ Appendix C Additional Implementation Details
For competing methods, we adopt the hyperparameters from their official implementations for VideoDirector wang2025videodirector, Light-A-Video zhou2025light, RelightVid fang2025relightvid, and Cosmos-T1 alhaija2025cosmos. We replace base models of VidToMe li2024vidtome and Slicedit cohen2024slicedit with IC-Light zhang2025scaling, and therefore we align their classifier-free guidance scale and diffusion sampling steps with those in zhang2025scaling. Additionally, we set VidToMe’s local and global token‐merging ratios to 0.6 and 0.5, respectively, mirroring the setting of our approach. For Slicedit, we adjust the weighting factor $\gamma$ in Eq. (2) from the default 0.2 to 0.05 to better balance temporal coherence and instruction adherence. All other hyperparameters remain at their default values. The modified VidToMe serves as the baseline of our model design.
During implementation, each $\kappa(x,y,t)$ comprises three components: (1) a per-pixel flow ID, (2) a quantized RGB color, and, optionally, (3) a world-frame voxel coordinate. For (1), flow IDs are derived from the optical flow estimated by the state-of-the-art MemFlow method dong2024memflow and the binary mask obtained by thresholding the soft mask in Eq. (6) of the main paper (values > 0.5 are set to 1; otherwise 0). In the initial frame, pixels receive unique flow IDs from 0 to HW1HW-1, where $H$ and $W$ denote image height and width. In subsequent frames, a pixel inherits the flow ID of its predecessor if connected by an unmasked flow; otherwise, it is assigned a new ID. This injects motion priors into the UVT representation. For (2), we quantize RGB values to 7 bits, ensuring that all pixels sharing the same UVT element differ by less than $2/255$ in any channel. This constraint mitigates erroneous flows that escape the mask and reinforces representation in regions exhibiting view-dependent effects. For (3), when per‐frame depth maps are available, they are reprojected into a point cloud using the camera intrinsics and extrinsics to determine world‐frame coordinates. This point cloud is then voxelized at a specified voxel size, and each pixel’s voxel coordinate is appended to $\kappa(x,y,t)$, yielding a more compact representation of static regions. For the CARLA dosovitskiy2017carla and InteriorNet InteriorNet18 datasets, voxel sizes are set to 0.05 m and 0.02 m, respectively. Notably, dynamic objects at different timestamps may spatially overlap in 3D, but they remain distinguishable by their flow IDs and quantized RGB colors. Consequently, each object at each timestep is represented by a distinct set of UVT elements, while the $\mathcal{L}_{1}$ temporal consistency loss preserves object identity across frames.
@@ Appendix D User Study
We conducted an online user study with 78 anonymous participants, evaluating 19 randomly selected video–text pairs from our datasets. The compared methods were IC-Light* zhang2025scaling, VidToMe li2024vidtome, Slicedit cohen2024slicedit, Cosmos-Transfer1 alhaija2025cosmos, Ours-light, and Ours-full. A screenshot of the questionnaire interface is shown in Fig.˜7. For each question, methods were anonymized and relighted videos were presented in random order; participants selected the two most preferred results. In compliance with the NeurIPS Code of Ethics, each participant received a compensation of \$0.70. Besides, we ensured that all collected data remained confidential and was not disclosed to any institutions or individuals.
Since each video spanned 10–20 seconds, completing the questionnaire took on average 13.5 minutes. Submissions requiring less than four minutes were deemed unreliable and excluded, yielding 65 valid responses. Fig.˜8 reports the frequency with which each method was chosen among the top two. Our full model achieved the highest preference rate, while the light variant ranked second. Although IC-Light* and VidToMe follow instructions well (cf. Tab. 2 of the main paper), their inferior temporal consistency make them much less preferred by users. Finally, we computed Bradley–Terry preference scores bradley1952rank as a comprehensive metric of user preference, as presented in the Tab. 2 of the main paper.
@@
Appendix E Physical Plausibility
This section illustrates how our model maintains physical plausibility. The physical plausibility of our method is inherited from IC-Light (zhang2025scaling), which is pre-trained on high-quality Light Stage data and has learned a physically grounded relighting process. Our main contribution lies in improving temporal consistency without altering the illumination priors embedded in the base model.
As detailed in Sec.˜3.2, we introduce a video model inflation mechanism based on token merging/unmerging and decayed multi-axis denoising to enable temporal feature-level information exchange. Since these components do not alter the prior knowledge encoded in the base model, the distribution of edited illumination aligns with that of IC-Light while enhancing temporal consistency.
In Section Sec.˜3.3, we propose a two-stage post-optimization strategy. The first stage smooths global exposure transitions using an appearance embedding, following practices in physically-based rendering methods like NeRF-W (mildenhall2021nerf) and 3DGS (kerbl3Dgaussians). The second stage refines local fluctuations without altering the overall lighting. Thus, our final results maintain the physically plausible qualities of IC-Light, while significantly improving temporal coherence. As shown in Fig.˜3 of the main paper and the video results, our illumination remains qualitatively aligned with IC-Light and VidToMe, but with fewer artifacts and greater temporal stability.
Thanks to the strong priors of IC-Light, our method focuses on temporal coherence and computation efficiency. Compared to optimization-heavy approaches such as Nerfactor (zhang2021nerfactor) and InvRender (zhang2022invrender), our post-optimization stage takes only around 2 minutes, with the entire pipeline completing in  10 minutes—substantially faster than training NeRF or 3DGS models, as discussed in Sec.˜2.2 of the main paper
Appendix F Social Impact